A monopolist operates under two plants, A and B. The marginal costs of these plants are given by 360 — 14x — 2x2and 310 — 15x — x2 with x representing units of output produced by each plant. If the price of the product is given by 396 — 4x2 , calculate the overall marginal cost and determine profit maximizing output in each plant.


1
Expert's answer
2020-10-12T03:39:03-0400

Plant 1: Mc1 "=360-14x-2x^2"

Plant 2: Mc2 "=310 - 15x -x^2"


C1 = "\intop"Mc1 dx = "360x-7x-2x^3\/3"


C2 = "\int"Mc2 dx = "310x-15x^2\/2-x^3\/3"


Total lost of the monopolist "=" C1 +C2

"360x-7x^2-2x^3\/3+310x-15x^2\/2-x^3\/3"


"670x-7x^2-15x^2\/2-2x^3\/3-x^3\/3"


"670x-(14x^2-15x^2)\/2-x^3"


"670x-(29x^2)\/2-x^3"


Overall MC "= 670-29x-3x^2"

For monopolist profit will be max.

"MR=MC"

"TR=P\u00d7Q"

"TR=(396-4x^2)x=396x-4x^3"

"MR=dTR\/dx=396-12x^2"


in Plant 1 profit will be max

"MR=MC"1

"396-12x^2=360-14x-2x^2"

"36=-14x+10x^2"

hence

"-10x^2+14x+36=0"

"10x^2-14x-36=0"

"5x^2-7x-18=0"

"x=2.722"

hence "x=3" units


in Plant 2

"MR=MC"2

"396-12x^2=310-15x-x^2"

"86=-15x+11x^2"

"-11x^2+15x+86=0"

"11x^2-15x-86=0"

"x=3.5598"

hence "x=4" units